OpenStudy (anonymous):

intweval noatation for x<1/3 and x ≥ 10/3

OpenStudy (anonymous):

there is no such thing. should be OR

OpenStudy (anonymous):

\[(-\infty,\frac{1}{3})\cup[\frac{10}{3},\infty)\]